• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Zheng

    Log entry time 15:00:23 on February17,2009

    Entry number 261640

    keyword=MCC update again

    There is beam "scrapping" problem going beyond 20uA while we need to go up to 30uA for the BCM calibration. The spokesperson is giving MCC 1 hour to fix the problem and we will re-evaluate the situation by 4pm if not fixed.